OpenStudy (anonymous):

Abby works in the sales department. This month the company is offering a 3-week paid vacation to every employee who sells 230% more this month than last month. If Abby sold $48,000 products last month, how much does she need to sell this month to earn the vacation?

OpenStudy (anonymous):

HELLLLLLLLLLLLLLP

OpenStudy (anonymous):

You guys are seriously life savers at this moment im terrible at math but get a `s all the time in reading

OpenStudy (anonymous):

so $48,000 = 100% we need x = 230% using ratios, ($48,000/100%)=(x/230%) when we cross multiply we get 100x=[230($48,000)] divide both sides by 100: x=[230($48,000)]/100 x=110,400

OpenStudy (mathteacher1729):

We can set up a ratio like this: \[\frac{x}{48,000}=\frac{230}{100}\] The left hand side says "x over 48,000.." and the right hand side says ... "is 230 percent". We want the value of x, so cross-multiply and solve. :) Stairway has the correct figure, though you should put a dollar sign to indicate units. Hope this makes sense!